### v1

# The infrared luminosity of retired and post-starburst galaxies: A cautionary tale for star formation rate measurements

Abstract
In galaxies with significant ongoing star formation there is an impressively tight correlation between total infrared luminosity ($L_{\mathrm{TIR}}$) and H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$), when H $\alpha$ is properly corrected for stellar absorption and dust attenuation. This long-standing result gives confidence that both measurements provide accurate estimates of a galaxy’s star formation rate (SFR), despite their differing origins. To test the extent to which this holds in galaxies with lower specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is the stellar mass), we combine optical spectroscopy from the Sloan Digital Sky Survey (SDSS) with multi-wavelength (FUV to FIR) photometric observations from the Galaxy And Mass Assembly survey (GAMA). We find that $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ increases steadily with decreasing H $\alpha$ equivalent width ($W_{\mathrm{H}\,\alpha}$, a proxy for sSFR), indicating that both luminosities cannot provide a valid measurement of SFR in galaxies below the canonical star-forming sequence. For both ‘retired galaxies’ and ‘post-starburst galaxies’, $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ can be up to a factor of 30 larger than for star-forming galaxies. The smooth change in $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, irrespective of star formation history, ionisation or heating source, dust temperature or other properties, suggests that the value of $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ is given by the balance between star-forming regions and ambient interstellar medium contributing to both $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$. While $L_{\mathrm{H}\,\alpha}$ can only be used to estimate the SFR for galaxies with $W_{\mathrm{H}\,\alpha}>3$ ($\mathrm{sSFR}\gtrsim 10^{-11.5}$/yr), we argue that the mid- and far-infrared can only be used to estimate the SFR of galaxies on the star-forming sequence, and in particular only for galaxies with $W_{\mathrm{H}\,\alpha}>10$ ($\mathrm{sSFR}\gtrsim 10^{-10.5}$/yr). We find no evidence for dust obscured star-formation in post-starburst galaxies.
keywords:
galaxies: evolution – galaxies: star formation – galaxies: starburst – galaxies: stellar content.
1 Introduction
Measuring the rate at which galaxies form stars is one of the key tools used to understand how they form and evolve. The field has benefited from a wide range of independent methods, from data spanning the full electro-magnetic spectrum, allowing for internal consistency checks and validation (see Kennicutt & Evans, 2012, for a review). This has led to significant confidence amongst extra-galactic astronomers in our ability to accurately measure the star formation rate (SFR) of any galaxy, of any type, at any redshift, with important caveats often under-appreciated.
Here we focus on one particular relation, between total infrared (TIR) luminosity (including mid- and far-IR) and H $\alpha$ luminosity. Once H $\alpha$ has been properly corrected for stellar photospheric Balmer absorption and for dust attenuation via the Balmer decrement method, the two luminosities are impressively tightly correlated in local star-forming galaxies (Rosa-González et al., 2002; Kewley et al., 2002; Rosario et al., 2016), despite their different physical origins. Balmer emission lines are produced via recombination of electrons to the $n=2$ level of the hydrogen atom; in star-forming galaxies this predominantly occurs following photoionisation and subsequent recombination and cascade within the H ii regions that surround hot (O and B type) stars. Mid to far-IR emission is caused by the heating of dust grains through the absorption of ultra-violet (UV) to optical photons. In highly star-forming galaxies these photons come predominantly from young, hot stars and interact with the dense, dust rich birth clouds of the star forming regions, leading to the strong relation between TIR emission and SFR. Unfortunately, neither of these SFR estimators are free of contamination or complications.
The basic physics of Balmer emission in star-forming galaxies is simple and well understood, with only photons with energies greater than 13.6 eV (Lyman continuum photons) able to contribute to the ionisation of the hydrogen atoms in the first place, meaning O and B stars are the primary contributor to Balmer emission in star-forming galaxies. However, there are two key complications in converting a H $\alpha$ emission line luminosity into a SFR. These are the attenuation of the photons as they travel through the galaxy ISM once leaving the H ii regions, and the contribution to the far-UV interstellar radiation field (ISRF) from active galactic nuclei (AGN), shocks and hot low-mass evolved stars (HOLMES; Flores-Fajardo et al., 2011), such as hot post-asymptotic giant branch stars and white dwarfs (Stasińska et al., 2008; Cid Fernandes et al., 2011). Through modelling the stellar continuum in SDSS galaxies using population synthesis models, Cid Fernandes et al. (2011) showed that galaxies with a low equivalent width of the H $\alpha$ line ($W_{\mathrm{H}\,\alpha}<3$) had an ionisation source consistent with the HOLMES present in the galaxy continuum spectrum, calling them ‘retired galaxies’. In these galaxies $L_{\mathrm{H}\,\alpha}$ is not proportional to their SFR, but is proportional to the total stellar mass of lower mass stars available to provide the ionising photons, as well as the presence of a reservoir of warm gas (see Herpich et al., 2018, for a comparison of retired galaxies with and without H $\alpha$ emission).
Once the H $\alpha$ photon is emitted from the H ii region close to the site of star formation, it can be absorbed and scattered by dust within the galaxy before being received by our telescopes [^2] We are ignoring here the possibility that dust grains within the H ii region might absorb the photon before it ever reaches the hydrogen atoms to ionise them. Such competition for Lyman continuum photons is difficult to detect (e.g. Charlot & Longhetti, 2001; Murphy et al., 2011).. In the case that some of the line emission still escapes and the dust distribution in the galaxy is smooth, the loss of light can be corrected for via the known ratio between Balmer line fluxes, combined with a knowledge of the dust attenuation as a function of wavelength (the Balmer decrement method, see e.g. Groves et al., 2012). In the case of clumpy dust the detected Balmer emission is weighted towards lower dust regions, and therefore the dust attenuation in the galaxy is underestimated using this method (Vale Asari et al., 2020). Comparison between spatially resolved and integrated light observations indicate that this effect is at the $\lesssim 10$% level in the majority of galaxies (Vale Asari et al., 2020; Belfiore et al., 2023), however the relative lack of very high spatial resolution integral field spectroscopy makes this estimate uncertain. At the extreme end of this scenario, there is clearly no way to correct for Balmer emission that has been entirely blocked by dense dust clouds.
On the other hand, the physics of mid- to far-IR luminosity is highly complex, with dependence on the gas density, dust-star geometry, the distribution of dust grain sizes and types which can vary with metallicity, and the shape of the ISRF, which depends on the stellar population balance (e.g. Draine & Li, 2007; Paladini et al., 2007; Bernard et al., 2010; Nersesian et al., 2019). It is predominantly the strong correlation with other measurements that provides the evidence for the link between TIR and SFR in high SFR galaxies. While Balmer emission lines are somewhat impacted by HOLMES, significant dust heating is caused by all stars with masses below O and B type stars. Diffuse dust emission permeates the entire galaxy contributing significantly to the IR luminosity of galaxies or regions with lower ongoing SFRs (e.g. Helou, 1986; Lonsdale Persson & Helou, 1987; Paladini et al., 2007; Bendo et al., 2010; Kennicutt et al., 2011; Calzetti, 2013; Nersesian et al., 2019). As a result, the TIR to H $\alpha$ luminosity ratio ($L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$) has been observed to increase with decreasing specific SFR ($\mathrm{sSFR}=\mathrm{SFR}/M_{*}$, where $M_{*}$ is a galaxy’s stellar mass) in integrated SDSS–Herschel comparisons (Rosario et al., 2016) and spatially resolved PHANGS–MUSE comparisons (Belfiore et al., 2023). While Rosario et al. (2016) put the effect down to selection bias, Belfiore et al. (2023) interpreted it as due to contamination of the IR luminosity by diffuse ISM dust.
The fact that there may be significant reservoirs of dust invisible to optical astronomers due to the enormous optical depths of stellar birth clouds has led to the idea that FIR may in fact be a better indicator of SFR in certain types of galaxies. This has become a critical discussion in particular in relation to galaxies which are currently shutting down their star formation (so-called ‘quenching’), and has crucial importance as the field grapples with understanding the wide range of processes that cause galaxies to stop forming stars. If the dust in a galaxy is sufficiently dense to bury light from a significant number of the O and B stars, as well as their surrounding H ii regions, then we have the case of a ‘dust obscured starburst’. Such a galaxy may reveal its presence in the optical via strong Balmer absorption lines, as the longer living A and F stars emerge from the dense clouds shrouding the younger O and B stars (Smail et al., 1999; Poggianti & Wu, 2000; Geach et al., 2009; Baron et al., 2022).
However, an excess Balmer absorption line strength is exactly the signature used to detect galaxies which have recently and rapidly shut down their star formation. Ironically, it is in these rapidly quenching galaxies where we might also expect to see the largest fraction of IR dust emission coming from the ambient ISM, unrelated to current star formation, as clearly demonstrated using hydrodynamical merger simulations by Hayward et al. (2014). This is because (a) we expect the diffuse dust emission located far from previous sites of star formation to still be present, (b) there is significant UV–optical light emitted by the A and F stars, (c) these A and F stars may also still be co-located with the remnant dust from the starburst episode. So while H $\alpha$ disappears rapidly following the starburst, total IR luminosity is expected to remain high, leading to an order of magnitude over-estimation in star formation, if using IR luminosity, from the point at which star formation starts to decline, right into the quenched phase many 100’s of Myr later (see figures 2 and 3 of Hayward et al. 2014).
As the debate about the role of total IR emission as a SFR indicator in quenching galaxies continues, and the origin of IR emission in ‘retired’ galaxies has not yet been studied, this paper takes a fresh look at the relationship between total IR luminosity and H $\alpha$ luminosity in a wide range of galaxies, including star-forming, post-starburst and retired galaxies. In Section 2 we present the catalogues that we have combined in our analysis, in Section 3 we summarise the relevant methods, and in Section 4 we present our results. We discuss our results with the help of a simple toy model in Section 5 and summarise in Section 6. Throughout the paper we assume a cosmology of $\Omega_{M}=0.3$, $\Omega_{\Lambda}=0.7$ and $h=0.7$ when converting fluxes into luminosity. All derived catalogues that we use assume a Chabrier (2003) initial mass function for conversion of stellar light into stellar masses. Unless otherwise stated $L_{\mathrm{H}\,\alpha}$ refers to dust attenuation and aperture corrected H $\alpha$ luminosity throughout the paper.
2 Data and sample selection
Our analysis requires a large sample of galaxies with good quality optical continuum spectroscopy to provide continuum subtracted, dust attenuation corrected H $\alpha$ luminosity, combined with aperture-matched UV, optical and mid- to far-IR multiwavelength photometric observations to provide accurate dust luminosities. To this end, we combine the spectroscopic galaxy sample from the Sloan Digital Sky Survey (SDSS) Data Release 7 (DR7, Abazajian et al., 2009), with photometry from the Galaxy And Mass Assembly (GAMA) Data Release 4 equatorial survey (Driver et al., 2022).
From SDSS we take nebular emission line fluxes, errors and equivalent widths from the starlight–SDSS catalogue (Cid Fernandes et al., 2009; see also Mateus et al., 2006; Stasińska et al., 2006; Asari et al., 2007; Cid Fernandes et al., 2010). The underlying stellar continuum is subtracted prior to measurement of the emission line fluxes using the starlight full spectral fitting package (Cid Fernandes et al., 2005). Stellar masses are calculated from the mass-to-light ratio of the starlight model, and corrected for aperture loss using the fiber-to-total $z$-band luminosity fraction. The luminosity of the H $\alpha$ emission line is corrected for dust attenuation using the Cardelli et al. (1989) law with $R_{V}=3.1$ and assuming an intrinsic Balmer decrement of $\mathrm{H}\,\alpha/\mathrm{H}\,\beta=2.86$.
From GAMA we obtain 21 band multi-wavelength photometry from the lambdar catalogue (Wright et al., 2016), which is aperture-matched and deblended, with variations in the point-spread function and pixel scale across the various bands properly accounted for (see also Driver et al., 2016). Most importantly for this paper, the GAMA equatorial fields include observations from Data Release 1 of the Herschel–ATLAS far-IR space telescope survey (H-ATLAS, Valiante et al., 2016; Bourne et al., 2016), and the Wide Field Infrared Survey Explorer which covers the mid-IR (WISE, Wright et al., 2010; Cluver et al., 2014). These two surveys provide 6 band photometry across the crucial 22–500 µm wavelength range333i.e. Wise W4, PACS 100, PACS 160, SPIRE 250, SPIRE 350, SPIRE 500. required to measure and characterise the emission from dust in galaxies. The majority of the galaxies in our sample also have GALEX NUV and FUV fluxes detected at $>3\sigma$, including the post-starburst and retired galaxies. The latter are typically detected with lower significance. We make use of the GAMA DR4 magphys catalogue, which takes input photometry from the lambdar catalogue using the code of da Cunha et al. (2008) to measure stellar population and dust properties for GAMA galaxies.
From the SDSS spectroscopic main galaxy sample (primtarget$=64$, Strauss et al., 2002) we exclude a small number of objects for which the SDSS 1D spectroscopic pipeline has identified problems, i.e. if the z_warning flag is set, or the velocity dispersion is fitted to the maximum value of v_disp$=850$ kms-1 (Stoughton et al., 2002). We use topcat (Taylor, 2005) to match the SDSS and GAMA galaxies, using the catalogue right ascension and declination values and requiring an on-sky separation of $\arcsec<0.5\arcsec$. This results in a parent sample of 12,738 galaxies.
We note that the GAMA magphys catalogue contains dust emission parameters even for galaxies with no mid- or far-IR photometric detections; these are predictions based on the magphys model priors. We therefore restrict the sample to only those galaxies with $\geq 2$ photometric bands between 22–500 µm detected at $>3\sigma$ in the lambdar catalogue. This reduces the sample to 4,368 galaxies. Finally, we restrict our sample to objects with H $\alpha$ detected at $>5\sigma$, leading to a final sample of 4,121 objects.
2.1 Post-starburst sample
There are many ways to identify post-starburst galaxies, although all methods are based around identifying galaxies with an excess of A/F stars compared to their O/B star population, indicating a recent and rapid shut-off in star formation. Traditionally a complete absence of nebular emission lines was used to provide evidence for a lack of O/B stars (e.g. Dressler & Gunn, 1983; Couch & Sharples, 1987; Zabludoff et al., 1996; Goto et al., 2003). However, this method excludes the many interesting post-starburst galaxies with residual ongoing star formation and other ionisation sources. Therefore, most modern methods use an alternative method to determine a deficit of O/B stars relative to the A/F star population, such as information from the stellar continuum shape, the equivalent width of one of the emission lines, likely ionisation sources of the emission lines or longer wavelength photometry (Wild et al., 2007; Yan et al., 2009; Alatalo et al., 2016; Yesuf et al., 2017).
Due to our requirement for significant H $\alpha$ emission line flux, as well as mid- and far-IR photometric detections, we note that the post-starburst sample studied in this paper is biased to include only post-starburst galaxies with a significant interstellar medium, as well as sources to heat and ionise it, such as residual star-formation, shocks, HOLMES or AGN. However, as it is predominantly these galaxies which are argued to be dusty starburst interlopers in quenching samples (e.g. Poggianti & Wu, 2000; Baron et al., 2022), they are the relevant objects for this analysis.
Table 1: Position, catalogue identifiers and basic properties of the post-starburst galaxies studied in this paper.
RA Dec Plate MJD Fiberid GAMA-DR4 Catalogues $z$ $M_{*}$ $L_{\mathrm{H}\,\alpha}$ $L_{\mathrm{TIR}}$ $W_{\mathrm{H}\,\alpha}$
CATAID dex(M) dex(L) dex(L) Å
130.77146 1.14878 467 51901 397 300757 5 0.078 10.4 7.6 10.8 4.1
133.40527 1.71763 469 51913 338 323854 2 0.058 10.7 8.3 10.4 7.7
134.17632 1.03737 468 51912 557 372076 2 0.088 10.7 7.4 9.9 2.9
135.19133 -0.49369 470 51929 290 550327 2 0.041 9.5 7.6 10.1 24.0
135.50417 -0.56457 470 51929 251 550389 2 0.089 9.9 7.6 10.4 14.6
135.88658 1.21010 470 51929 402 372446 2,4,5 0.058 9.8 6.1 9.5 0.6
136.75280 0.06094 470 51929 115 210106 2,5 0.164 10.4 7.8 10.6 4.8
137.33520 0.73873 470 51929 621 623161 2 0.054 10.2 7.6 9.8 12.5
139.67269 2.02517 473 51929 464 383051 2 0.087 11.0 8.2 10.7 6.3
140.82538 0.56146 474 52000 408 217011 2 0.088 10.4 7.9 9.9 3.6
140.90214 2.13637 473 51929 593 383312 2 0.017 10.6 7.5 10.3 1.9
140.92896 1.34917 473 51929 29 373374 2 0.025 11.0 6.7 10.0 1.2
177.64780 -2.63937 329 52056 117 123917 3 0.093 10.6 7.9 10.5 13.3
177.65639 1.50016 515 52051 295 397294 1,2,4,5 0.078 10.2 6.5 9.9 1.0
178.10642 -1.26744 330 52370 361 143889 1,2,5 0.060 10.4 6.5 10.1 2.3
179.98362 -0.68304 285 51663 162 40161 4 0.080 10.3 7.6 10.1 10.6
180.74100 0.05999 286 51999 395 70638 2,4 0.080 10.9 8.0 10.5 9.0
181.11809 -0.88879 286 51999 208 536202 2 0.094 10.0 8.4 10.4 43.5
181.22794 0.31118 286 51999 477 610664 1 0.100 10.7 7.8 10.6 5.1
181.43242 1.07895 286 51999 444 23386 4 0.086 10.7 7.8 10.5 9.2
181.62258 -0.55774 286 51999 150 560745 2 0.095 10.6 8.1 10.6 5.4
182.43333 -0.28872 286 51999 78 55486 2 0.160 10.2 8.6 11.0 54.5
182.70929 0.47087 286 51999 634 85480 4 0.078 10.4 7.9 10.3 13.5
183.03908 -1.68483 332 52367 527 138549 2,5 0.103 10.2 6.4 9.8 0.7
212.55941 -0.57859 302 51688 110 567624 3 0.026 9.6 6.9 9.2 14.2
213.29913 -0.39941 304 51609 291 62531 3 0.126 10.8 8.1 10.7 14.6
214.16447 1.30599 533 51994 55 227568 2 0.123 10.2 8.2 10.3 17.8
214.19902 -1.07059 303 51615 89 37000 2,3,5 0.125 11.1 8.7 10.6 2.6
214.46337 -0.13896 304 51609 216 592879 4 0.169 10.8 8.3 10.6 14.1
216.59955 2.26955 535 51999 341 342615 2,5 0.105 10.7 8.2 10.6 12.6
216.81288 0.93103 306 51637 321 106634 2 0.052 10.4 7.8 10.1 17.4
221.69920 -1.21974 920 52411 460 493601 2,5 0.083 10.4 8.3 10.4 5.1
223.07965 0.13545 308 51662 638 79784 2 0.147 10.6 7.9 10.8 18.1
Catalogue references: (1) Goto (2007) (2) Wild et al. (2007); (3) Alatalo et al. (2016);
(4) Pattarakijwanich et al. (2016); (5) C. Tremonti and Chen et al. (2019).
We match our combined SDSS+GAMA+H-ATLAS/WISE sample onto several SDSS DR7 post-starburst samples in the literature, identifying 33 post-starburst galaxies in total. Positions, catalogue identifiers and basic properties for these galaxies are given in Table 1. The sample includes 3 galaxies from Goto (2007), 24 from Wild et al. (2007) [^4] We follow Pawlik et al. (2018) and exclude dusty PCA-selected post-starburst candidates with H $\alpha$/H $\beta$ ratios $>6.6$. These are possible contaminants for this particular method, due to the impact of significant dust attenuation on the shape of the stellar continuum., 4 from Alatalo et al. (2016), 7 from Pattarakijwanich et al. (2016) and 9 applying the method described in Chen et al. (2019) to SDSS DR7, referred to as the ‘Tremonti sample’. Some objects are identified using multiple methods, with the largest cross over being 8 objects contained in both the Wild et al. (2007) and Tremonti samples. The selection criteria of each of these samples is summarised in Table 2.
Table 2: Comparison between the different selection methods of post-starburst galaxies included in this paper. In order to distinguish absorption from emission in this table, equivalent width values ($W$) follow the convention of positive for absorption and negative for emission. Note that in the remainder of the paper we use absolute equivalent width values for $W_{\mathrm{H}\,\alpha}$ in emission.
Author Survey Balmer absorption lower limit Ongoing SFR upper limit
Goto (2007) SDSS DR5 $W_{\mathrm{H}\,\delta}>5$ $W_{\mathrm{H}\,\alpha}>-3$ and $W_{[\mathrm{O}\,\textsc{ii}]}>-2.5$ [^5] Lines measured with flux summing method (Goto et al., 2003).
[…]
Alatalo et al. (2016) SDSS DR7 $W_{\mathrm{H}\,\delta}>5$ Line ratios consistent with shock ionisation and inconsistent with star-formation ionisation.
Pattarakijwanich et al. (2016) SDSS DR9777Sample restricted to DR7 only in this paper. Template fitting [^8] Method from Quintero et al. (2004). $\mathrm{A/(A+K)}>0.25$ and $W_{\mathrm{H}\,\delta}>4$ $W_{[\mathrm{O}\,\textsc{ii}]}>-2.5$
Tremonti sample; Chen et al. (2019) SDSS DR7 $\mathrm{H}\,\delta_{\rm A}>3$ [^9] Lick Index (Worthey & Ottaviani, 1997). $W_{\mathrm{H}\,\alpha}>-10$ and $\log_{10}|W_{\mathrm{H}\,\alpha}|<0.23\times\mathrm{H}\,\delta_{\rm A}-0.46$
The sample includes 7 post-starburst galaxies where optical emission lines indicate that the dominant ionising source is residual star-formation (below the demarcation line of Stasińska et al. 2006), 8 are consistent with being ionised by HOLMES ($W_{\mathrm{H}\,\alpha}<3$), 6 ionised by AGN (above the demarcation line of Kewley et al. 2001 and $W_{\mathrm{H}\,\alpha}>3$), and 10 lie between the Stasińska et al. (2006) and Kewley et al. (2001) demarcation lines of the [N ii]/H $\alpha$ and [O iii]/H $\beta$ line ratios diagram (Baldwin et al., 1981). For the purposes of this paper it is essential to include galaxies with all ionisation sources, we therefore refrain from converting $L_{\mathrm{H}\,\alpha}$ into SFR for our figures, as in many cases this would be incorrect.
3 Methods
Here we summarise the relevant data analysis and interpretation methods used in this paper.
Partial correlation coefficients: Throughout the results section we use partial correlation coefficients to quantify the correlation between three variables. For variables ($x$, $y$ and $z$) we firstly compute the Spearman rank-order correlation coefficient between each: $\rho_{xy}$, $\rho_{xz}$ and $\rho_{yz}$. The partial correlation between $x$ and $z$, given $y$, is then:
$\rho_{xz|y}=\frac{\rho_{xz}-\rho_{xy}\rho_{zy}}{\sqrt{(1-\rho_{xy}^{2})}\sqrt{(1-\rho_{zy}^{2})}}.$ (1)
The angle of the correlation with $z$, projected onto the $x$$y$ plane ($\theta_{p}$), is given by:
$\tan{\theta_{p}}=\frac{\rho_{xz|y}}{\rho_{yz|x}}.$ (2)
An angle of $45^{\circ}$ indicates the $z$ quantity correlates equally with the $x$ and $y$ quantities; $0^{\circ}$ indicates the $z$ quantity correlates with $y$ alone; $90^{\circ}$ indicates the $z$ quantity correlates with $x$ alone; $-45^{\circ}$ indicates the $z$ quantity correlates equally with the $x$ and with $y$, but an increase in $x$ corresponds to a decrease in $y$.
Expected luminosity ratio: We compare our observed ratio between the luminosities of H $\alpha$ and total IR (TIR), to the ratio expected if both quantities measure SFR. Following the notation of Kennicutt & Evans (2012), $\log_{10}(\mathrm{SFR})=\log_{10}(L_{X})-\log_{10}(C_{X})$, with SFR in M/yr and luminosities in erg/s. With $C=43.41$ for TIR (3–1100 µm) and $C=41.27$ for H $\alpha$ (Hao et al., 2011; Murphy et al., 2011), we obtain the expected relation for star-forming galaxies of:
$\log_{10}(L_{\rm TIR})=\log_{10}(L_{\mathrm{H}\,\alpha})+2.14.$ (3)
We note that these conversions have a number of assumptions underlying them, most notably of constant SFR for 10 Myr for $L_{\mathrm{H}\,\alpha}$ and 100 Myr for $L_{\mathrm{TIR}}$. We will discuss the implications of this approximation at length in Section 5.
magphys fitting: Although the magphys catalogue was computed by the GAMA team and made available as part of their data release, it is important to understand the underlying dust emission model in order to interpret our results. magphys is an energy balance code that assumes that all stellar light absorbed from the UV–optical wavelength range is re-emitted in the mid- and far-IR. It fits a two component dust emission model, with one contribution from dust in the stellar birth clouds, and the second from dust in the ambient ISM (see figure 2 in da Cunha et al. 2008). Both components contribute to flux emitted by polycyclic aromatic hydrocarbons (PAHs) and hot mid-IR continuum. The birth clouds additionally contribute warm dust in thermal equilibrium, while the ambient ISM contributes cold thermal dust. Free parameters are the fraction of total IR luminosity contributed by dust in the ambient ISM, the temperature of the two thermal components, and relative fractions of light contributed by the different birth cloud components.
We note that the total IR luminosity output by magphys is calculated from the energy lost in the UV–optical wavelength regime, which is then distributed via the model between 3–1000 µm. Although $L_{\mathrm{TIR}}$ is necessarily model dependent due to the sparse sampling of the IR SED by observations, comparisons between methods indicate that this does not introduce significant errors (e.g. Hunt et al., 2019).
Modified black body fitting: To ensure that our results are not dependent on the magphys model, we independently fit the total IR luminosity with a modified black body, using the python MCMC package emcee (Foreman-Mackey et al., 2013). We assume a fixed dust emissivity index $\beta=1.8$ and a dust mass absorption coefficient $\kappa_{850}=0.077\mathrm{\,m^{2}\,kg^{-1}}$ and include a mid-infrared power-law to account for hot dust, allowed to vary uniformly between $5.50.5<\alpha<$ 5.5 (see also Casey, 2012; Bourne et al., 2019). We fit for a temperature range between 10 and 100 K using a uniform prior. We obtain a very tight 1:1 correlation between our measurement of $L_{\rm TIR}$ and the value from magphys, with a Spearman rank correlation coefficient of 0.99 when all 6 mid- to far-IR bands are detected at $>3\sigma$, decreasing slightly to 0.96 when only 3 bands are detected.
[…]
4 Results
Figure 1: The total dust (mid-far infrared) luminosity versus H $\alpha$ luminosity for galaxies observed in both SDSS and GAMA/Herschel-ATLAS/WISE (small dots), where $L_{\mathrm{H}\,\alpha}$ has been both dust and aperture corrected. Post-starburst galaxies are shown as large circles with the same colour scale. We only include galaxies with H $\alpha$ flux detected at $>5\sigma$ and fluxes in 2 or more 22–500 µm bands detected at $>3\sigma$. The arrow and angle indicate the direction of partial correlation with the $z$ (colour) axis. The line indicates equity in the case that both luminosities trace the galaxy SFR, assuming standard conversions from the literature. No smoothing has been applied to the data. Left: all galaxies, regardless of the origin of the H $\alpha$ flux. Right: only galaxies where the H $\alpha$ flux comes predominantly from star-formation, as determined from their emission line ratios (see text for details).
The left panel of Figure 1 presents the total dust luminosity ($L_{\mathrm{TIR}}$), as estimated by magphys between 3–1000 µm versus the dust and aperture corrected H $\alpha$ luminosity ($L_{\mathrm{H}\,\alpha}$) for all galaxies in our sample, colour coded by the equivalent width of H $\alpha$ ($W_{\mathrm{H}\,\alpha}$). We immediately see that the expected correlation between $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ only holds for galaxies with $W_{\mathrm{H}\,\alpha}\gtrsim 3$Å, the limit below which a galaxy is classified as ‘retired’ (Cid Fernandes et al., 2011). The arrow indicates the direction of partial correlation ($\theta_{p}=-56^{\circ}$), confirming $W_{\mathrm{H}\,\alpha}$ to be strongly positively correlated with $L_{\mathrm{H}\,\alpha}$, and negatively with $L_{\mathrm{TIR}}$. As we will see from the next plot, the stronger correlation with $L_{\mathrm{H}\,\alpha}$ arises purely due to the presence of the retired galaxies.
The right hand panel of Figure 1 restricts the sample to the 1,638 galaxies with optical emission line ratios that indicate the gas is predominantly ionised by young stellar populations, using the criteria of Stasińska et al. (2006) which is based on the commonly used [N ii]/H $\alpha$ and [O iii]/H $\beta$ line ratios. It is noteworthy that there is still a trend with $W_{\mathrm{H}\,\alpha}$, with the conversion line equating both luminosities to SFR only holding for galaxies with $W_{\mathrm{H}\,\alpha}\gtrsim 10$ (Hao et al., 2011; Murphy et al., 2011; Kennicutt & Evans, 2012). The partial correlation angle of $\theta_{p}=-48^{\circ}$ indicates that $W_{\mathrm{H}\,\alpha}$ traces $L_{\mathrm{TIR}}$ just as much as it does $L_{\mathrm{H}\,\alpha}$ and we can see that accounting for $W_{\mathrm{H}\,\alpha}$ would account for most of the scatter in the relation. The results shown in these figures are unchanged if we use a simple modified black body fit to the mid- and far-infrared data (see Section 3).
In Figure 1 the post-starburst galaxies are indicated by larger circles. The identical colour scaling of the post-starburst galaxies relative to the background population of galaxies demonstrates that post-starburst galaxies are not unusual in their dust luminosities, given their H $\alpha$ equivalent widths. The majority of the post-starburst galaxies in this sample have $W_{\mathrm{H}\,\alpha}>3$, and in Figure 1 they lie on the upper side of the star-forming sequence in $L_{\mathrm{TIR}}$ versus $L_{\mathrm{H}\,\alpha}$. These are interpreted as being in the process of shutting down their star formation, or have alternative ionisation sources. A small number have $W_{\mathrm{H}\,\alpha}\lesssim 3$, i.e. consistent with being entirely quenched, and the majority of the line emission arising from older stars present in the galaxy. These are the objects with the highest excess $L_{\mathrm{TIR}}$.
Our results show that all post-starburst galaxies have dust luminosities larger than predicted by $L_{\mathrm{H}\,\alpha}$, but consistent with star-forming galaxies or retired galaxies with the same $W_{\mathrm{H}\,\alpha}$. If these post-starburst galaxies are in fact ‘dust obscured starbursts’, then the excess $L_{\mathrm{TIR}}$ common to all galaxies with $W_{\mathrm{H}\,\alpha}<10$ should be interpreted as due to complete obscuration of the H $\alpha$ line. In the following sections we will discuss why we do not think this is the case.
Figure 2: Total galaxy stellar mass versus H $\alpha$ luminosity for the same galaxy samples as in Figure 1. H $\alpha$ luminosity has been corrected for dust attenuation and aperture, as described in the text. Dashed lines indicate constant sSFR as labelled, in units of yr-1. Arrows and angles indicate the direction of partial correlation with the $z$-axis quantity. No smoothing has been applied to the data. Top left: coloured by total infrared luminosity divided by dust and aperture corrected H $\alpha$ luminosity. The expected log(ratio) is 2.14 based on standard conversions from luminosity to SFR. Top right: coloured by total infrared luminosity. Bottom left: coloured by fraction of the total dust luminosity coming from the ‘diffuse’ ISM, as fit by magphys. Bottom right: coloured by the temperature of the cool ‘diffuse’ ISM dust component, as fit by magphys. In this panel we require galaxies to have $>5\sigma$ detections in one additional dust band compared to our main analysis.
To further investigate the cause of the varying $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio with $W_{\mathrm{H}\,\alpha}$, Figure 2 shows the same samples plotted as stellar mass $M_{*}$ versus $L_{\mathrm{H}\,\alpha}$. The top left panel largely repeats the information presented in Figure 1, but the inclusion of stellar mass makes the blue and red sequences stand out more clearly. Interestingly, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio changes smoothly across this figure, with no obvious discontinuity between the blue and red sequences, or for the post-starburst galaxies. The partial correlation angle of $\theta_{p}=-38^{\circ}$ is again consistent with $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ trending most closely with sSFR. The smoothness of the transition suggests that whatever is responsible for the varying ratio has no knowledge of the origin of the emission, or whether the galaxy is ‘quenched’, recently quenched or still star-forming.
Red sequence galaxies clearly have the highest $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$, with the majority having a ratio more than 1 dex greater than the expected ratio based on standard conversions (2.14, see Section 3), while 19 galaxies have a ratio $\geq 1.5$ dex greater. At the other extreme, galaxies with either very high $L_{\mathrm{H}\,\alpha}$/$M_{*}$ (equivalent to sSFR for galaxies where $L_{\mathrm{H}\,\alpha}$ is dominated by star-formation) or very high $L_{\mathrm{H}\,\alpha}$ have a ratio that is lower than expected, with 32 having a ratio $\geq 0.5$ dex lower. Again we see that post-starburst galaxies, whether completely quenched or not, have the expected $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio compared to other galaxies with the same stellar mass and $L_{\mathrm{H}\,\alpha}$.
The top right panel colours the galaxies by $L_{\mathrm{TIR}}$. This visualisation adds no new data, but again demonstrates the continuity between the blue and red sequence. $L_{\mathrm{TIR}}$ depends almost equally on $L_{\mathrm{H}\,\alpha}$ and $M_{*}$, i.e. at fixed $L_{\mathrm{H}\,\alpha}$, $L_{\mathrm{TIR}}$ increases with $M_{*}$, again showing that both quantities cannot trace SFR. The smoothness of the colour scaling across the blue and red sequence again shows that $L_{\mathrm{TIR}}$ is entirely agnostic as to the origin of the emission or quenched status of the galaxy.
The lower left panel colours the galaxies by the fraction of total dust luminosity that is fit by the ambient ISM dust component in magphys ($f_{\mu,\mathrm{IR}}$), as opposed to the birth cloud dust surrounding the star-forming regions. Here we see a more distinct discontinuity between the blue and red sequences, with the majority of red sequence galaxies requiring $\lesssim 10$% additional contribution from a hotter birth cloud component, while the blue sequence galaxies have typically $\gtrsim 50$% contribution from hotter birth clouds.
The lower right panel colours the galaxies by the temperature of the ambient ISM dust component fit by magphys. Constraining dust temperature accurately requires more dust bands than a simple dust luminosity, we therefore required galaxies to be detected at $5\sigma$ in 3 or more 22–500 µm bands for this analysis (2,877 galaxies). While there is a significant scatter, the predominant trend is for galaxies with higher $L_{\mathrm{H}\,\alpha}$ to have higher temperature ISM dust ($\theta_{p}=-15^{\circ}$). There is a small residual trend with stellar mass, but not to the same extent as for $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. This shows that the changing ISM dust temperature is not the cause of the changing luminosity ratio. magphys also fits for the temperature of the birth cloud thermal component (not shown). We find a similar result, with galaxies with higher $L_{\mathrm{H}\,\alpha}$ having higher temperature birth cloud dust and any residual trend with stellar mass being small.
We investigated various other correlations in the dataset, which are not shown for conciseness, but we describe a couple more here. Dust attenuation, as measured by both the emission line Balmer decrement and by the magphys multi-wavelength photometric fit, follows the same trend as $L_{\mathrm{TIR}}$ in blue sequence galaxies i.e. higher mass, higher $L_{\mathrm{H}\,\alpha}$ galaxies have greater attenuation. However, the trend does not extend to the red sequence, where there is lower attenuation overall ($\tau_{\rm V,Balmer}\lesssim 0.5$). The relative luminosity fraction of the warm and cold thermal components fit by magphys are correlated predominantly with stellar mass, with a small residual trend with $L_{\mathrm{H}\,\alpha}$. High mass galaxies and red sequence galaxies have significantly higher fraction of dust fit with a cold thermal component than lower mass, blue sequence galaxies. This provides a consistent picture to that shown for $f_{\mu,\mathrm{IR}}$, indicating ambient ISM dust heating depends on stellar mass via the amount of UV and optical photons available from lower mass stars, and therefore contributes most significantly to $L_{\mathrm{TIR}}$ in red sequence galaxies.
4.1 Post-starburst galaxies
[…]
5 Discussion
In the standard picture (e.g. Helou, 1986; Lonsdale Persson & Helou, 1987; Draine & Li, 2007), a galaxy ISM contains a temporally stable ambient reservoir of dust that is built up steadily via many generations of stars. Alongside this ambient dust, there are temporally transient dense birth clouds surrounding new star-forming regions; the greater the number of star-forming regions, the greater the number of dense birth clouds. When these young star-forming regions dominate over the ambient dust, we find that $L_{\mathrm{TIR}}$ correlates strongly with $L_{\mathrm{H}\,\alpha}$ and both provide a good estimate of SFR in galaxies. As the specific SFR of the galaxy declines, $L_{\mathrm{TIR}}$ becomes increasingly dominated by the ambient dust. This standard picture of dust in galaxies is the basis of the two component dust model of Charlot & Fall (2000), which is the model implemented in magphys, as well as some other SED modelling codes (e.g. GRASIL, Domínguez-Tenreiro et al., 2014).
In fact, the sources of $L_{\mathrm{H}\,\alpha}$ in galaxies can be thought of in much the same way as the sources of $L_{\mathrm{TIR}}$: a background population of hot, low mass, evolved stars (HOLMES) provides the temporally stable equivalent of the ambient dust that builds up over time, while the temporally transient hot OB stars provide the majority source of Lyman continuum photons when present. Not entirely coincidentally of course, the timescale during which we expect OB stars to stop providing Lyman continuum photons is of the same magnitude as the timescale that we expect the dense birth clouds to disperse. Therefore, as sSFR decreases, the stellar population moves from OB star dominated to G+ star dominated, at the same rate that the dust clouds move from birth cloud dominated to diffuse ISM dominated.
However, $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ respond differently to the stellar population mix. Because $L_{\mathrm{H}\,\alpha}$ emission caused by HOLMES is very weak, $L_{\mathrm{H}\,\alpha}$ drops very sharply once the star-forming regions vanish. On the other hand, $L_{\mathrm{TIR}}$ remains high because dust can be heated by all stars emitting in the UV and optical wavelength regime. Thus, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio increases, reflecting the higher fraction of dust luminosity coming from ambient ISM compared to the fraction of $L_{\mathrm{H}\,\alpha}$ coming from older stars. We hypothesise that it is this that causes the smooth trend between $L_{\mathrm{TIR}}$, $L_{\mathrm{H}\,\alpha}$ and $M_{*}$ continuing into the red sequence, without discontinuity, despite the apparently different ionising and heating mechanisms, and star formation histories of the galaxies. The changing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ with sSFR then simply reflects the differing relative importance of OB stars to HOLMES for $L_{\mathrm{H}\,\alpha}$, and diffuse ISM to birth clouds for $L_{\mathrm{TIR}}$ emission.
As discussed in Section 3, when luminosities are converted into a SFR, a constant SFR over a fixed timescale is assumed (10 Myr for $L_{\mathrm{H}\,\alpha}$ and 100 Myr for $L_{\mathrm{TIR}}$). While this might be a reasonable approximation for a star-forming galaxy, it clearly is not for a young post-starburst galaxy where the SFR has dropped precipitously over the past hundred Myr. To untangle the relative importance of star formation history versus assumed star-formation timescale to the excess $L_{\mathrm{TIR}}$ seen in retired, post-starburst and star-forming galaxies with $W_{\mathrm{H}\,\alpha}<10$, we turn to toy models.
Figure 3: Total infrared versus H $\alpha$ luminosity for the star-forming (SF), post-starburst (PSB) and retired toy models in $M_{*}$ bins. Triangles, squares and stars are used respectively for the middle, second-to-last and largest mass bins. The two lowest-mass bins for star-forming galaxies are shown as dots. Larger symbols indicate larger effective dust attenuation, with $\tau_{V}=0.1,0.5,1.0$ and $2.0$. The grey line is the SFR conversion line as in Figure 1.
5.1 Toy SED models
To illustrate how $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ depend on a galaxy’s stellar populations, we constructed a set of toy SED models based on average star formation histories for star-forming, post-starburst and retired galaxies in the starlight–SDSS main galaxy sample catalogue. We note that the post-starburst galaxies were selected via the Goto (2007) method, meaning they have completely ceased star-formation and thus match the subset of our post-starburst galaxies with $W_{\mathrm{H}\,\alpha}<3$. The models were constructed in mass bins of 1 dex using Bruzual & Charlot (2003) models with a Chabrier (2003) initial mass function. Further details are provided in Appendix A.
We calculated $L_{\mathrm{H}\,\alpha}$ from the extrapolated SED models in the UV by integrating the rate of ionizing photons, assuming the case B hydrogen recombination and an ionisation-bounded scenario (as in Stasińska et al., 2006; Cid Fernandes et al., 2011). For quenched post-starburst and retired galaxies, since the measured light fraction from young stars is over-estimated due to limitations in the stellar population models, we only considered stellar populations older than $10^{8}$ yr (see Ocvirk, 2010; Cid Fernandes et al., 2011, equation 2 and footnote 6).
From the intrinsic model spectrum $L_{\lambda}^{\mathrm{int}}$, we modelled a dust reddened spectrum assuming a single dust screen,
$L_{\lambda}^{\mathrm{obs}}=L_{\lambda}^{\mathrm{int}}e^{-\tau_{V}q_{\lambda}},$ (4)
where $\tau_{V}$ is the effective dust attenuation in the $V$-band and $q_{\lambda}\equiv\tau_{\lambda}/\tau_{V}$ is the shape of the dust attenuation law, here taken to be the Cardelli et al. (1989) extinction curve with $R_{V}=3.1$. We created models with four values of the effective dust attenuation, $\tau_{V}=0.1,0.5,1.0$ and $2.0$. $L_{\mathrm{TIR}}$ was then calculated by integrating the difference between the intrinsic and dust reddened spectrum, and assuming all light attenuated by dust is re-radiated in the mid- and far-IR [^10] Changing the attenuation law to a Charlot & Fall (2000) two-component dust law using $q_{\lambda}=(\lambda/{5500\,\mathrm{\AA}})^{-0.7}$, and for a birth cloud $V$-band attenuation twice that of the ISM, results in a modest increase in the predicted $\log L_{\mathrm{TIR}}$ of 0.16 dex for the SF and of 0.08 dex for the post-starburst and retired galaxies..
[…]
When going from models for star-forming galaxies to post-starburst/retired galaxies, within the same mass and $\tau_{V}$ bin (fixed symbol style and size, changing colour) there is a slight decrease in $L_{\mathrm{TIR}}$ as old stars still heat the diffuse ISM, and a huge drop in $L_{\mathrm{H}\,\alpha}$. This highlights the fact that HOLMES are a much less powerful ionising source than OB stars: the rate of ionising photons decreases by five orders of magnitude from a 10 Myr to a 1–10 Gyr stellar population; see figure 1, panel b of Cid Fernandes et al. (2011).
The slightly higher $L_{\mathrm{TIR}}$ of the quenched post-starburst galaxies compared to retired galaxies is due to the presence of a large number of relatively hot A and F stars to heat the dust, rather than the recent precipitous drop in SFR (which from Figure 4 can be seen to be insignificant for well over the 100 Myr averaging time assumed for $L_{\mathrm{TIR}}$ in star-forming galaxies). The comparatively small difference between post-starburst and retired galaxies, compared to the difference between post-starburst and star-forming galaxies, shows that the precise stellar population balance is a much less important factor than the current specific star formation rate in determining the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio.
Our toy models do not include post-starburst galaxies that are in the process of quenching, which are seen in Figure 1 to lie just above the star-forming sequence. Here we expect a combination of a rapidly declining recent SFR and a reduction in birth cloud dust to be contributing to the enhanced $L_{\mathrm{TIR}}$ compared to $L_{\mathrm{H}\,\alpha}$ (see also Hayward et al., 2014).
5.2 Dust-star geometry, interstellar radiation field and cloud density
Interestingly, the many complications that could arise in terms of dust temperature, dust-star geometry, dust cloud density and even ionisation source do not appear to be important to our results, at least to first order, even for post-starburst galaxies which should provide a test of the most extreme quenching conditions. At far-IR wavelengths, the temperature of the dust is thought to be determined mostly by the dust-star geometry, rather than the shape of the interstellar radiation field (ISRF, e.g. Bernard et al., 2010). Thus, for any given stellar population mix, the stars which emit the most UV–optical photons will contribute most to the total $L_{\mathrm{TIR}}$, i.e. OB stars in star-forming galaxies, AF stars in post-starburst galaxies, or G and later in retired galaxies, but the differing temperature of the stars is not relevant. Our results in Figure 2 suggest that dust temperature predominantly depends on $L_{\mathrm{H}\,\alpha}$, with a small residual mass trend. Galaxies with the highest $L_{\mathrm{H}\,\alpha}$ in our sample have higher ambient ISM and dense BC dust temperatures, and this is expected to be due to the closer proximity of the dust to the star-forming regions, rather than the changing ISRF (e.g. Nersesian et al., 2019).
The lack of any appreciable difference in the temperature of post-starburst galaxy ISM dust compared to normal galaxies is interesting (bottom right panel of Figure 2), as we might expect the A/F stars to be co-located with the majority of the ISM dust following the end of the starburst, and therefore able to enhance the heating in a similar way to the brightest starbursts. Rowlands et al. (2015) found a decrease in post-starburst dust temperature with time since burst. However, they were studying generally younger post-starburst galaxies than in this sample, and temperature was estimated from colour index rather than the magphys fit as used here. It seems likely that the dust heating observed in the youngest post-starbursts in their sample was due to a changing balance of hot birth cloud dust to cooler ISM dust following the star-formation episode, rather than a more general heating of the ambient ISM i.e. closer to the $f_{\mu,\mathrm{IR}}$ parameter in the bottom left panel of Figure 2.
5.3 SFR estimators and full spectral fitting
The debate about the importance, or otherwise, of dust heating by lower mass stars has a long history, with both sides likely ‘right’ in their context. In early works, the inability to correct $L_{\mathrm{H}\,\alpha}$ for stellar absorption and dust attenuation led to the conclusion that IR was a better total SFR indicator for galaxies, regardless of what was heating it (e.g. Lonsdale & Helou, 1985). However, this argument became obsolete once better optical data became available (e.g. Rosa-González et al., 2002; Kewley et al., 2002) and the two were shown to be exceedingly tightly correlated in local star-forming galaxies.
Another argument is that it does not matter where the photons come from, they still represent star formation, and should be counted even if the star formation happened long ago. This is because combining $L_{\mathrm{TIR}}$ with a UV luminosity to account for the unextincted photons accounts for most photons generated by stars that are still alive at the time of observation, regardless of their mass or when they were formed. However, this argument is only plausibly useful when integrating in both time and space, not when trying to estimate the ongoing SFR of a particular galaxy, or set of galaxies as it does not tell us the number of stars recently formed.
A more important, yet subtle, discussion is around the timescales for different SFR estimators. To obtain the current SFR of a galaxy, the basic underlying principle is to count the number of stars just formed. This is, however, not easy for spatially unresolved data – in reality we have to count the brightest, and therefore hottest, stars formed, then assume an initial mass function to give us the SFR. Unfortunately, even this is hard to do, so we resort to proxy estimators that are necessarily time averaged.
For $L_{\mathrm{H}\,\alpha}$ the time-averaging is the main sequence lifetime of O and B stars, about 10 Myr. It is this short timescale that makes it attractive as a robust SFR indicator. For FIR, the time averaging is around 100 Myr for star-forming galaxies, which accounts for lower mass stars also heating the dust (Kennicutt, 1998). All conversions between a proxy luminosity and SFR require the assumption of a constant SFR over the averaging timeframe. For some galaxies, such as those which are currently rapidly declining in SFR, the longer estimators are therefore less suitable. Our results, supplemented by our toy models, highlight the subtle difference between star formation history versus contamination by ambient ISM dust in enhancing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. The post-starburst galaxies in our sample that are currently decreasing in SFR, but still have appreciable SFR, have a slightly enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, and they lie close to ordinary star-forming galaxies. This could be due to their rapidly declining star formation history, as well as their reducing contribution of birth cloud dust to $L_{\mathrm{TIR}}$. On the other hand, the post-starburst galaxies that have completely quenched their SFR have a very substantially enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, causing them to lie close to the retired galaxies. When the observed luminosity is dominated by photons from significantly lower mass stellar populations, we are no longer measuring the number of recently formed stars, but rather the integrated mass formed during the entire lifetime of the galaxy. This leads to a strong correlation between $L_{\mathrm{TIR}}$ and $M_{*}$, rather than $L_{\mathrm{TIR}}$ and SFR. This suggests that we should see different partial correlation angles between $L_{\mathrm{TIR}}$, $M_{*}$ and $L_{\mathrm{H}\,\alpha}$ for star-forming, retired and post-starburst galaxies. Unfortunately the small numbers of retired and post-starburst galaxies in our samples made it impossible to tease these relations apart, and a larger set of retired and quenched post-starburst galaxies would be required.
In principle, obtaining an SFR from full multi-wavelength SED fitting, rather than simple proxy estimators, should solve these problems (Hunt et al., 2019). However, this requires the underlying modelling assumptions to represent reality, both in terms of the star formation histories and the dust emission models. Hunt et al. (2019) shows that different SED fitting models and proxy indicators disagree the most when sSFR is low, suggesting that not all SED fitting models include the correct assumptions (or model ‘priors’) in this regime.
5.4 Comparison to previous work
Although papers studying the SFR calibrators commonly investigate galaxy types, few provide measurements of the $W_{\mathrm{H}\,\alpha}$ or sSFR of the galaxies being used for calibration. One exception is Hao et al. (2011) where the majority of galaxies have $W_{\mathrm{H}\,\alpha}>10$, and their figure 7 shows an overestimation of the SFR using an FIR-based calibrator by $0.3$ dex for the small number of galaxies with $W_{\mathrm{H}\,\alpha}>5$. This is entirely consistent with our results.
Rosario et al. (2016) looked at $L_{\mathrm{H}\,\alpha}$ versus $L_{\mathrm{TIR}}$ in an early Herschel-SDSS sample similar to that studied here, and while they did not link their results to star formation histories, their plots show the same excess of $L_{\mathrm{TIR}}$ in low sSFR galaxies. It is reassuring that an entirely different analysis leads to the same quantitative results.
Our results are also consistent with those of Nersesian et al. (2019) using the SED modelling code CIGALE (Boquien et al., 2019), who noted that galaxies with $\log_{10}\mathrm{(sSFR/yr}^{-1})>-10.5$ had dust mainly heated by radiation emitted by the young population, whereas lower sSFR had dust heated by the older population. This corresponds to $W_{\mathrm{H}\,\alpha}>10$, matching our results.
And finally, our results are consistent with the spatially resolved PHANGS–MUSE SFR calibration comparisons of Belfiore et al. (2023), who noted a systematic increase in IR-based SFR estimators compared to dust attenuation corrected $L_{\mathrm{H}\,\alpha}$ at low specific star formation rates of $\log_{10}\mathrm{(sSFR/yr}^{-1})<-10$.
All these results show that IR-based SFR indicators should only be used where $W_{\mathrm{H}\,\alpha}>10$ or $\log_{10}\mathrm{(sSFR/yr}^{-1})>-10.5$.
6 Summary
We have compared the H $\alpha$ luminosity, corrected for dust and aperture effects, to the total IR luminosity of 4,121 galaxies observed in both spectroscopic SDSS DR7 and GAMA DR4 catalogues, with detections in the mid- and far-IR observations from WISE and Herschel–ATLAS. We find that:
  • $L_{\mathrm{TIR}}$ is strongly correlated with $L_{\mathrm{H}\,\alpha}$ for galaxies with $W_{\mathrm{H}\,\alpha}$ $\gtrsim 10$Å or $\log_{10}\mathrm{(sSFR/yr^{-1})}>-10.5$, with a ratio that is consistent with locally determined calibrations. In this regime, both $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ can be used to estimate integrated SFRs of galaxies.
  • However, for galaxies with $W_{\mathrm{H}\,\alpha}\lesssim 10$ or $\log_{10}\mathrm{(sSFR/yr^{-1})}<-10.5$, $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ increases by up to 1.5 dex. In this regime, $L_{\mathrm{TIR}}$ should not be used to measure the integrated SFRs of galaxies. We remind readers that $L_{\mathrm{H}\,\alpha}$ can only be used to estimate the SFR of galaxies when $W_{\mathrm{H}\,\alpha}\gtrsim 3$ (Cid Fernandes et al., 2011).
  • The smoothness in the relationships between $L_{\mathrm{TIR}}$, $L_{\mathrm{H}\,\alpha}$ and galaxy stellar mass for galaxies, regardless of their dominant heating or ionisation source, is consistent with the ‘standard’ 2-component picture for dust emission in galaxies. A background of old, low mass stars both heat the dust and provide a low level background of ionising photons in all galaxies, regardless of sSFR. As sSFR increases, the number of star-forming regions increases, and these then dominate the heating and ionisation budget.
  • The reason for the increasing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio with decreasing sSFR is simply due to the fact that dust is heated by UV–optical photons, while $L_{\mathrm{H}\,\alpha}$ relies on Lyman continuum photons alone. Thus, $L_{\mathrm{H}\,\alpha}$ decays rapidly once star formation ceases with old stars providing only a very small number of far-UV photons, while dust continues to emit, presumably until the grains are destroyed.
  • We use toy models to show that the excess $L_{\mathrm{TIR}}$ is predominantly due to the increasing dominance of old, low mass stars heating the dust, and not due to recent star formation history. However, the rare population of galaxies with ongoing, yet rapidly declining, star formation rates (young post-starburst galaxies) may have slightly enhanced $L_{\mathrm{TIR}}$ simply due to the longer averaging time of $L_{\mathrm{TIR}}$ than $L_{\mathrm{H}\,\alpha}$ when converted into an SFR.
  • Recently and rapidly quenched post-starburst galaxies do not show any evidence of hosting dust-obscured starbursts. Their high $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratios are entirely consistent with being caused by the above processes.
An improved understanding of heating and ionisation mechanisms in retired and post-starburst galaxies could be gained from spatially resolved observations, both in the mid- and far-IR and with optical integral field spectroscopy. Sensitive radio observations would also provide additional assurance that significant amounts of dust obscured star formation was not being missed from the optical census of the Universe. However, our results indicate that the high fraction of rapidly quenched galaxies detected in optical surveys is not a result of high levels of contamination from dust-obscured starbursts, and therefore understanding rapid quenching pathways to quiescence remains of strong importance to understanding galaxy evolution.
Data Availability
The SDSS DR7 data is available from http://classic.sdss.org/dr7/. The GAMA DR4 catalogues are available from https://www.gama-survey.org. starlight catalogues are available on request to NVA or VW.
Acknowledgements
We would like to thank Rob Kennicutt and Desika Narayanan for their help and advice in interpreting the results, and Kenneth Duncan, Ho-Seong Hwang and Leah Morabito for sharing data and information on datasets during preliminary investigations carried out prior to this paper being written.
[…]
GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalogue is based on data taken from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programmes including GALEX MIS, VST KiDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT and ASKAP providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO, and the participating institutions. The GAMA website is https://www.gama-survey.org/.
Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS web site is www.sdss.org.
SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, the Chilean Participation Group, the French Participation Group, Harvard-Smithsonian Center for Astrophysics, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
Appendix A Star formation histories for toy SED models
To create our toy SED models we selected galaxies from the starlight–SDSS catalogue which belong to the SDSS main galaxy sample, with a signal-to-noise ratio $S/N\geq 10$ at 4020 Å, $M_{*}\geq 10^{7}$ M and non-negative Petrosian half-light radius, which results in a parent sample of 616,097 objects. The star-forming sample comprises 133,769 galaxies below the Stasińska et al. (2006) line and with $W_{\mathrm{H}\,\alpha}\geq 3\,\mathrm{\AA}$. Our subsample of 314,880 retired galaxies is defined as those with $W_{\mathrm{H}\,\alpha}<3\,\mathrm{\AA}$, while the 580 post-starburst were selected using the Goto (2007) criteria.
Figure 4 shows the average starlight fitted star formation histories of our SF, post-starburst and retired galaxy subsamples divided into five stellar mass bins (Cid Fernandes et al., 2005). These average star formation histories are used to calculate $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ in our toy models, as described in Section 5. We excluded the two lowest mass bins for the retired and post-starburst samples, due to the low number of galaxies or very low predicted luminosities. The visible bump at 1 Gyr lookback time in the star-forming galaxy star formation histories, and to a lesser extent in the retired galaxy star formation histories, is caused by imperfections in the stellar population models. The youngest age bin is similarly affected and is removed from the analysis, as described in Section 5 (Ocvirk, 2010; Cid Fernandes et al., 2011). The logarithmic time axis is scaled deliberately to highlight the recent past history of these galaxies, which is the important part for our analysis. The post-starburst galaxies have had a notably stronger burst of star formation between 500 and 1.5 Gyr in lookback time than caused by the model imperfections, but have experienced little star formation since then. The quiescent galaxies have been quiescent for at least 1 Gyr.
These star formation histories show that the star formation rate of all of these populations has been very stable for the last 100 Myr. This is important for our discussion of the cause of the excess $L_{\mathrm{TIR}}$ in the retired and post-starburst galaxies in Section 5.
Figure 4: Average star formation histories for the retired, post-starburst (PSB) and star-forming (SF) subsamples in stellar mass bins, used in the toy models in Section 5. The amount of mass formed as a function of stellar population age $t_{\star}$ is represented by the initial mass-fraction population vector $\mu_{\mathrm{ini}}$.